OpenStudy (anonymous):

Estimate the integral using the midpoints of 100 intervals. fnInt(1/sqrt(a+x^2))dx {0,1}

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(1)\div \sqrt{1+x ^{2}}dx\]

OpenStudy (anonymous):

let x = tanu ...

OpenStudy (anonymous):

I am stuck as to how i would set up a sequencee equation.

OpenStudy (zarkon):

\[\int\limits_{a}^{b}f(x)dx\approx\sum_{i=1}^{n}f\left(a+\left(i-\frac{1}{2}\right)\frac{b-a}{n}\right)\frac{b-a}{n}\]

OpenStudy (zarkon):

let \(a=0,b=1,n=100\) and \[f(x)=\frac{1}{\sqrt{1+x^2}}\]

OpenStudy (anonymous):

ln|x+sqrt(x^2+1)|

OpenStudy (zarkon):

I get .88137358702

OpenStudy (anonymous):

yea thats actually the correct answer. I just didnt know the process in getting that answer. Thank you.

OpenStudy (zarkon):

yw